OpenStudy (anonymous):

Doug wants to buy apples and oranges. He can spend up to $5; apples cost $1 per pound and oranges cost $1.25 per pound. How can Doug spend the $5? If the solution to the problem is graphed, it will be an inequality with a line as the border. What does the line represent?

OpenStudy (igreen):

Your inequality is: \(1x + 1.25x \le 5\).

OpenStudy (igreen):

If your inequality has \(\le\) or \(\ge\) the line will be SOLID. If your inequality has < or > the line will be DOTTED.

OpenStudy (igreen):

\(\bbox[5pt, #000000 ,border:5px solid #00FF00]{\bf\huge \color{#00FF00}\sim \! \! \! \! \color{#00FF00}i\it\Huge\color{#00FF00}G\color{#00FF00}r\color{#00FF00}e\color{#00FF00}e\color{#00FF00}n\bf\color{#00FF00}ツ}\)

OpenStudy (anonymous):

thanksss